OpenStudy (mokeira):

in my country's education system, we were taught NOTHING on radians. I will be majoring in math soon and i need to master stuff with radians in it and it seem sooo hard to me. does someone care to explain to mean basics like what is a radian and any other useful info, please?

OpenStudy (rational):

http://i.imgur.com/WbJwn5J.gif

OpenStudy (anonymous):

>.< i dont remember any radian in univ though

OpenStudy (mokeira):

nice!!! @rational

OpenStudy (mokeira):

there is some stuff i need to have completed before we start calculus...and its the only part remaining @BSwan

OpenStudy (mokeira):

its like math basics generaly, not necesarily related to calculus @BSwan

OpenStudy (rational):

`radian` is simply a measuring unit like meter or foot or degrees

OpenStudy (mokeira):

how many degrees=1radian?

OpenStudy (rational):

like u measure distance in meters, u can measure angle in radians

OpenStudy (rational):

don't jump to the relationship between radians and degrees yet, first convince urself on how many radians a complete circle makes

OpenStudy (rational):

whats the formula for circumference of a circle ?

OpenStudy (mokeira):

so degrees and radians have no relation whatsover. I already got that part @rational but when getting are is where i am confused. when you have been given radians

OpenStudy (anonymous):

oh , today i had this argument , when i had calcules 1 i took A , cuz every thing seems to be repeated from high school , but now i just relalized that calcules 1 added nothing to my knowleged , instead it make all the rusty stuff in head , if time come back , i will study it on my own >.< so be carfull and be creative in it

OpenStudy (anonymous):

:O

OpenStudy (mokeira):

my bro told me the same thing..told me to jump to calculus II or some other thing...but i decided to just start fresh

OpenStudy (mokeira):

how do i find area when given radians @rational

OpenStudy (anonymous):

no no im not saying jump to calc 2 :O (big mistake ) also its not allowed in major sheet :P im saying to not rely on ur teacher :O be creative in studing it , also read more that one resource :O

OpenStudy (mokeira):

lmao @rational TRUE!!!!!!!

OpenStudy (rational):

you should get use to angle in radians, radians is more natural than degrees

OpenStudy (rational):

what is given in radians ? and what area u want to find ?

OpenStudy (mokeira):

@BSwan i had an option to do so coz aparently im too good in math....so said the college guys. we are using this aleks site...i dont realy get their explanations on that particular topic of radians

OpenStudy (mokeira):

is there a basic formula for finding area of a sector of a circle when given radians... @rational

OpenStudy (anonymous):

yeah do you know the circle area ?

OpenStudy (rational):

well you know that "circumference = 2pi radians" use it to find the area of circle

OpenStudy (rational):

an angle of "2pi radians" creates an area of "pi r^2" so how much area an angle of "\(\theta\)" would make ?

OpenStudy (mokeira):

ok...say for example radius if 7 and you have also been given \[3\pi radians\]

OpenStudy (anonymous):

for example :- 

OpenStudy (rational):

Area of sector of \(\theta ~radians\) : \[\huge \pi r^2 \times \dfrac{\theta }{2\pi}\]

OpenStudy (mokeira):

wont pi and pi cancel?

OpenStudy (rational):

they will cancel, but i like remembering above form because it helps you see clearly how you scaling the original area based on angle

OpenStudy (mokeira):

oooh, ok i guess i will jus have to do practice questions on the same then

OpenStudy (anonymous):